Required information

[The following information applies to the questions displayed below.]

You have just been hired as a financial analyst for Lydex Company, a manufacturer of safety helmets. Your boss has asked you to perform a comprehensive analysis of the company’s financial statements, including comparing Lydex’s performance to its major competitors. The company’s financial statements for the last two years are as follows:

Lydex Company
Comparative Balance Sheet
This Year Last Year
Assets
Current assets:
Cash $ 930,000 $ 1,170,000
Marketable securities 0 300,000
Accounts receivable, net 2,580,000 1,680,000
Inventory 3,570,000 2,200,000
Prepaid expenses 250,000 190,000
Total current assets 7,330,000 5,540,000
Plant and equipment, net 9,460,000 9,020,000
Total assets $ 16,790,000 $ 14,560,000
Liabilities and Stockholders' Equity
Liabilities:
Current liabilities $ 3,980,000 $ 2,920,000
Note payable, 10% 3,640,000 3,040,000
Total liabilities 7,620,000 5,960,000
Stockholders' equity:
Common stock, $70 par value 7,000,000 7,000,000
Retained earnings 2,170,000 1,600,000
Total stockholders' equity 9,170,000 8,600,000
Total liabilities and stockholders' equity $ 16,790,000 $ 14,560,000
Lydex Company
Comparative Income Statement and Reconciliation
This Year Last Year
Sales (all on account) $ 15,830,000 $ 13,280,000
Cost of goods sold 12,664,000 9,960,000
Gross margin 3,166,000 3,320,000
Selling and administrative expenses 1,402,000 1,592,000
Net operating income 1,764,000 1,728,000
Interest expense 364,000 304,000
Net income before taxes 1,400,000 1,424,000
Income taxes (30%) 420,000 427,200
Net income 980,000 996,800
Common dividends 410,000 498,400
Net income retained 570,000 498,400
Beginning retained earnings 1,600,000 1,101,600
Ending retained earnings $ 2,170,000 $ 1,600,000

To begin your assignment you gather the following financial data and ratios that are typical of companies in Lydex Company’s industry:

Current ratio 2.3
Acid-test ratio 1.1
Average collection period 40 days
Average sale period 60 days
Return on assets 8.9 %
Debt-to-equity ratio 0.7
Times interest earned ratio 5.9
Price-earnings ratio 10

3. You decide, finally, to assess the company’s liquidity and asset management. For both this year and last year, compute:

a. Working capital.

b. The current ratio. (Round your final answers to 2 decimal places.)

c. The acid-test ratio. (Round your final answers to 2 decimal places.)

d. The average collection period. (The accounts receivable at the beginning of last year totaled $1,640,000.) (Use 365 days in a year. Round your intermediate calculations and final answers to 2 decimal place.)

e. The average sale period. (The inventory at the beginning of last year totaled $2,000,000.) (Use 365 days in a year. Round your intermediate calculations and final answers to 2 decimal place.)

f. The operating cycle. (Round your intermediate calculations and final answers to 2 decimal place.)

g. The total asset turnover. (The total assets at the beginning of last year totaled $14,580,000.) (Round your final answers to 2 decimal places.)
Answer #1

Answer:

a. Working capital. = Current Assets - Current Liabilities

This Year = 7330000 - 3980000 = 3350000

Last Year = 5540000 - 2920000 = 2620000

b.Current Ratio = Current Assets / Current Liabilities

This Year = 7330000 / 3980000 = 1.84

Last Year = 5540000 / 2920000 = 1.897

c.Acid Test Ratio = Quick Assets / Current Liabilities

This Year = (930000 + 2580000 ) / 3980000 = 0.88

Last Year = (1170000 + 300000 + 1680000 ) / 2920000 = 1.08

d. The average collection period. = 365 / (Sales / Avg. Accounts Receivable)

This Year = 365 / ( 15830000 / ( 2580000 + 1680000 ) / 2 ) = 365 / (15830000 / 2130000 ) = 365 / 7.43 = 49.13 Days

Last Year = 365 / ( 13280000 / ( 1640000 + 1680000 ) / 2 ) = 365 / (13280000 / 1660000 ) = 365 / 8 = 45.63 Days

e. The average sale period = 365 / (Cost of goods sold / Avg. Inventory)

This Year = 365 / (12664000 / ( 3570000 + 2200000 ) / 2 ) = 365 / (12664000 / 2885000 ) = 365 / 4.39 = 83.14 Days

Last Year = 365 / ( 9960000 / ( 2000000 + 2200000 ) / 2 ) = 365 / (9960000 / 2100000 ) = 365 / 4.74 = 77 Days

f. The operating cycle = Average Collection Period + Average Sale Period

This Year = 49.13 Days + 83.14 Days = 132.27 Days

Last Year = 45.63 Days + 77 Days = 122.63 Days

g. The total asset turnover. = Sales / Avg. Total Assets

This Year = 15830000 / (16790000 + 14560000 ) / 2 = 15830000 / 15675000 = 1.009

Last Year = 13280000 / ( 14580000 + 14560000 ) / 2 = 13280000 / 14570000 = 0.911